# Ultra-Fast Language Generation via Discrete Diffusion Divergence Instruct

### v2

@@ F.2 Results on Zero-Shot Perplexities
F.3 Implementation Details of the Ablation Studies
@@
Score decompose. The baseline model is a one-step generator, mapping a fully masked input $\mathbf{z}_{t}$ ($t=1$) directly to the final output $\mathbf{x}$. With this technique, we decompose the generation into a two-step process $\mathbf{z}_{t}\rightarrow\mathbf{z}_{\tau}\rightarrow\mathbf{x}$ where $t=1$ and $\tau\in(0,1)$. The model first generates an intermediate state $\mathbf{z}_{\tau}$ at a randomly sampled time following $\pi(t)$ and then generates the final output $\mathbf{x}$ from $\mathbf{z}_{\tau}$. The ablation (‘w/o Score Decompose’) reverts to the direct, single-step generation.
@@
Coupled time $t$. This component synchronizes the time steps used for student generation and discriminator evaluation. In the standard setup, the intermediate denoising time $\tau$ for the student’s two-step generation ($\mathbf{z}_{t}\rightarrow\mathbf{z}_{\tau}$) and the corruption time $\tau^{\prime}$ for preparing the discriminator’s input ($\mathbf{x}\rightarrow\mathbf{z}_{\tau^{\prime}}$ for corrupted student generations, and $\mathbf{x}^{\prime}\rightarrow\mathbf{z}_{\tau^{\prime}}^{\prime}$ for corrupted teacher generations) are coupled, i.e., $\tau=\tau^{\prime}$. The ablation (‘w/o Coupled Time $t$’) decouples them by sampling $\tau$ and $\tau^{\prime}$ independently.
@@
Weight function $\omega(t)$ correction. The gradient of the objective (5) is weighted by a time-dependent factor $\omega(t)$. The ablation (‘w/o $\omega(t)$ Correction’) uses a constant weight, i.e., $\omega(t)=1$. Our full model applies a theoretically-motivated correction based on the noise schedule’s signal rate $\alpha_{t}$, setting $\omega(t)\propto-\alpha^{\prime}_{t}/(1-\alpha_{t})$, which re-weights the objective based on the difficulty of the denoising step at time $t$.
@@
Time scheduler $\pi(\cdot)$ weighting. This controls how $\tau$ is sampled and (optionally) whether we apply importance weighting, which also refers to the sampling distribution $\pi(\tau)$ for the intermediate time step $\tau$. The ablation setting (‘w/o $\pi(\tau)$ Weighting’) samples $\tau$ from a uniform distribution. Our method employs a non-uniform Beta distribution as $\pi(\tau)$ to focus the student’s training on specific, more informative intervals of the denoising trajectory. Specifically, we consider heavy-early ($\text{Beta}(2,5)$), and heavy-late ($\text{Beta}(5,2)$) schedules to improve sample quality. Practically, for smaller NFEs (e.g., 8 and 16 NFEs) we bias toward earlier times (e.g., $\text{Beta}(2,5)$); for larger NFEs (e.g., 32, 64, and 128 NFEs) we bias toward later times (e.g., $\text{Beta}(5,2)$).
@@
Regularization. To stabilize training and maintain generalization, we introduce two regularization terms to the student’s loss. The first is a KL divergence loss, which aligns the student’s output logits with those of the frozen teacher at both steps of the generation process ($\mathbf{z}_{t}\rightarrow\mathbf{z}_{\tau}$ and $\mathbf{z}_{\tau}\rightarrow\mathbf{x}$, where $t=1$ and $\tau\in(0,1)$). Among Forward/Backward/Jeffreys implementations, we found forward KL consistently yielded the most stable optimization and best validation in our cases, so we adopt it for all ablations and apply it with a weight of $0.05$ in the student objective. The second term is an entropy bonus on the student’s output distribution to encourage exploration, which is weighted by $0.0005$. The ablation (‘w/o Regularization’) removes both the KL-divergence and entropy terms from the student’s objective.
@@
Guided inference. As elaborated in Section 4.4, this technique is applied only during sampling to leverage the trained discriminator and improve sample quality. We employ a hybrid strategy that divides the denoising process into two phases. For the first 50% of NFEs, we use gradient tilting ($M=1$), where the guidance scale $h$ is linearly increased from 0 to 30 to steer the generation. For the remaining 50% of NFEs, we switch to multi-candidate re-ranking ($h=0$), sampling $M=4$ candidates at each step and selecting the one with the highest reward. The ablation (‘w/o Guided Inference’) disables this process ($h=0,M=1$), reverting to standard unguided ancestral sampling.
F.4 Ablation Study Result Analysis
@@
F.5 Model Scaling Up
@@
F.6 Protein Sequence Generation
Figure 5: pLDDT comparison between DiDi-Instruct and the DPLM-150M across different sequence lengths ($L=100,200,300,400,500$) and NFEs. Our method consistently outperforms the teacher, achieving up to +10 pLDDT gains at shorter sequence lengths (e.g., $L=100$) and maintaining superior structural confidence across all lengths, even with substantially fewer sampling steps. Moreover, the distilled student exhibits more stable performance across NFEs, whereas the teacher shows larger variability as the number of steps increases.
Dataset and evaluation metric. The UniRef50 dataset (Suzek et al., 2014) contains approximately 45 million protein sequences, comprising roughly 14 billion amino acid tokens. Following previous work (Wang et al., 2024), we adopt a vocabulary of 33 tokens and a maximum sequence length of 1024, with longer proteins chunked into shorter chunks. We compute the predicted local distance difference test (pLDDT) score using the ESMFold model (Lin et al., 2023) to assess the foldability of the generated protein sequences, with values above 70 indicating high structural confidence. We report pLDDT score across different sequence lengths and varying NFEs.
@@ F.6 Protein Sequence Generation
Figure 7: Visualization examples of high-confidence protein sequences ($\text{pLDDT}>70$) generated by the DiDi-Instruct model, with lengths ranging from 50 to 500 across different NFEs.
F.7 Text Examples
@@
Text obtained with $1024$ NFEs teacher model (1/3). Perplexity=$38.53$, Entropy=$5.22$. He was mid-morning early, and I started forming a motorcade for him. He stopped every NYPD car, and I radioed back to every bicycle officer that was in there, to protect him. I did not hear you said to be involved in it. That you were involved with that campaign. I never asked anybody to test it. You did need help you leaked it, not made it for yourself. He parked and it was then that I saw that president and ended up there, speaking with a journalist. The Prince was granted anonymity and noted that the FBI was aware of the familiarity with Russia and was conducting the investigation as appropriate. In some accounts, Casey Wilson, senior counsel for the Bonden Bisson, told The Post that the President is aware of the inquiry. But Mr. Bakers testimony is the first time someone has long taken something of concern into the hands of their officials probing Trumps personal relationships with foreign officials. Former FBI regional director James Kogan, a former officer who met with President Trump, was once briefed about the inquiry from the Russian District Attorney General. Victor Atkinson, an FBI chief-officer who is retired, hacked the personal PINs and cell phone lines of the two Trump officials to conceal the information, as shown inside the countrys electronic networks. Anonymous messages can appear in our apps like Twitter and Facebook and if ones political actions resists they become blocked. The Daily Signal columnist, and former Edward Snowden contractor, last summer. According to the piece he leaked, his company was set up to log phone calls for the officials so that some of Hillarys callers would end up because Trump and businesses required a gag. I was using this same setup last week during my private conversation with Trump. I called him after this piece of political reporting and "faming," he said of Barack Obama, as he told me last May. The conversation about the leak was livestreamed at 1 p.m. this evening. CAIRCLE MEDALE, RAC.Y. Authorities a 28-year-old Burnet County receiving a call at 10:30 p.m. live across the street in Greene County reported that the man he heard was telling a neighbor to kill him. Abedoy and Terry Angelo, 24, found the house. They were charged with disturbing the peace, said Greene Police Detective Nathaniel Clay. "There is a door with a phone in a glass covering, and there is a grenade at the window," Clay said. "The caller smashed out the window, indicating it was the mans father." Hes had a similar house for a week. They were just found. Police said Angelo said he thought up the name while he was delivering. "Police determined that the most recent break in the home had been set in place," Clay said. He said assistance.
Text obtained with $1024$ NFEs teacher model (2/3). Perplexity=$38.53$, Entropy=$5.22$. Montana State Supreme Court will vote Monday-2 to extend the reach of the case in Indiana until the league has their final ruling in court on November 4. The order in case reads as the first rule that the league will either settle trial court or the trial. With the lower court denying it. This means that when the ramifications stand, all players could be considered at "the game" like mayow or receive other incentives, likely a winning boost after Fridays game in Chicago. Gomez is 6-foot-6 by weight when he opens up. And probably is the future top free agent, but no players can return for taking a portion of those benefits. S. Thibodeau/AP Westwood is the standard locale to that end. Under the rules, the league must travel a minimum of four minutes per day to represent Finals 2 and 4. To every benefit from all the games, and that only players who participate play each others for each game. The league, which determines the future status of the highest free agent, will head to court in Westwood by Wednesday to avoid the headaches. The league serves as trial court in the case -- presumably with the court, which then gets to decide on any procedural cases that may arise -- before June 18, although the league says its not likely. The current clause disallows the NBA from linking any agent to the league. And prohibits it from accepting unany sponsorships that may have affected players without the approval of a director. Through its appeal, the NBA believes the Westwood will decide whether to receive any. It will update the league as it makes its decision the other day. If results are not appealed, the league ordered the court in it its ruling on March 29. It considered for an appeals court ruling on Monday, but that order has not been taken up by courts. The case seemed to already be dead Friday. After all of the courts first 18 officials handed the briefcase and could not earn a reach ruling in court itself, but it argued that it was hearing their appeal. The rule, when it is written, sends the 16-appointed back to a trial. And if they appeals the decision by following Friday to sea next Tuesday, hopefully the dispute will then be determined by the trial court. If the case is settled, the success of the Indiana incident in September could determine if the trial court can use its argument against the league and the NBA. Once both sides have signed a settlement agreement. The league says that under Indiana law the league must dismiss rulings Nov. 4 by the lower court, meaning that if it claims that a trial ruling will not be valid, it will be too late to qualify.
Text obtained with $1024$ NFEs teacher model (3/3). Perplexity=$38.53$, Entropy=$5.22$. And even less because to me the LGBT community are homophobic. Were really unforgressive because generally part of us is about equality. And I guess people really do not know, if were LGBT and were not going to pay for it, wouldnt we just do it? Well, yeah, yeah, yeah, that would be great. No, I mean realized we would pay for that because then youd just have a really bad time. If something is seriously important to you and wanting to be able to participate participate in it, youd be great at that. What do you think would be of stepping up to support your local team? Is there any way your parents would have to be able to if you played to the next level. Do they feel that responsibility as parents? Yes, but I do think that when you grow up in an LGBT environment, a lot of stuff is going to happen. Thats why it has been about anything that is civil. Look, I think that its working, but obviously a family can never get through that hierarchy here because younger people are very well theyre raised, and theyre going through a lot of bad experiences around nobody that are parents. Growing up from a parent figure is just a less enjoyable environment to be able to have. We are telling the story here, so you have an answer for it. How long have you and your band have toured a college district? When were you in college? Hunsheeling Foo since I was in college. Ive never really thought about recruiting for a high school. But it wasnt just just me going onto seeing whatever Neil did or why Google just made this about it all, or it was me going into every second week. Or going into at least some draft, though, maybe thats different than knowing what this thing is. Thats why I scholarship, that made the sense. I only broke records for two years at Kansas State and Kansas, I graduated and I had many great battles afterward, so theyre always in discussion. And I have a friend recruited as a part of the Michigan teams and Michigan. David Denselmoss, former Buffalo Bills backup QB for the Eagles that coach Bruce Allen used in recruiting from Michigan for Colorado. (He was originally planned to have been a Michigan transfer.) We happened to start your press conference because you mentioned high-profile recruits but not selected in that one, either. Dolting Peps: No question. Its certainly not an uncommon one to make that mistake. But if you do ask me theres nothing I can do wrong with being among the 60 worst coaches in football.
@@
Text obtained with 88 NFEs distilled by DiDi-Instruct (1/3). Perplexity=$62.24$, Entropy=$5.17$. An engineers run up against a system of dynamic memory-producing technology. An examiners mechanism of memory has to reduce property crimes against the technology. An engineers been up against a method of dynamic memory-producing thats used to reduce civilian property violence. Indias memory system has run into a deficit of memory and that economic damage in India has run up against a system of Now and Then. The new system will be designed as an economically secure and efficient system in determining how residents receive support. A loss of support from memory management systems wont go too far from promise. With a deficit of 1 million for residents, the Navajo Nation is testing a new system that wont prevent you from receiving your financial stream. The Navajo Nation faces a deficit of 1.88 million in systems that are designed to enhance what residents receive. Its a potential loss of support from a system that has been receiving both positive and negative feedback from the community. A presidential candidate who once served during four tours in Iraq and Afghanistan refers to his new record as a figurative use of pop culture. I got a disappointed order from the State Department, the new record is to go against the American health care system. This wasnt a response to his campaigns decision to use his new record for Donald Trumps campaign. I used a figurative term. The record is meant to go against the American health care system. I got a disappointed order from the State Department, no record for this. After decades of defending against the laughable claims of many affirmative-action groups, a study on the psychology of affirmative action could soon reveal the particular role of such programs at schools. The wake of the administrations efforts to eliminate affirmative action are endless. A memo released by the National Federation of Teachers found that threats exist for every affirmative action program back at the schools. In light of the Presidents efforts to address the seemingly endless claims, providers face other forms of administrative sanctions. The organizations campaign memo found that in the wake of the increased number of affirmative action programs, schools are at risk of diseases, fires and other issues. There are affirmative actions by particular providers which threaten every elementary and secondary school district.
Text obtained with 88 NFEs distilled by DiDi-Instruct (2/3). Perplexity=$62.24$, Entropy=$5.17$. A new website was created to accept yellow-brown channels in a continents Middle East country. The groups websites would like to be receiving the same TV channels as they block the spread of AIDS. This was a cause for concern. The task forces group had originally initiated a website to accept the AIDS TV channels, but accepting these channels would be bad for democracy, the group said. There was another website which was for the receiving countries of these channels. The GPAA task group also had to create a new site to accept yellow test channels into the continents same AIDS channels. At the time, the task force raised concerns about how accepting a TV channel would change the groups website. A report late this week said the app allows ride-drive interview subjects from education, billing-care, and at a cost of euros, chat to a user of politicians and smartphones to ask their questions. It will beg users to download the first cultural invasion app. The invasion also allows users to answer questions about their own methods, the IFP news agency said this week. The app allows ride-drive interviews with students at a cost of one million euros, but a user of politicians and smartphones can ask their questions. We have the moral obligation to ask, can we stop the way this government seeks to destroy our planet, culture and human rights? Jan Muir, director of the Swiss Research Council for Governance and Science, said hes not happy with what the U.S. is doing. This year a few established a new club in Pariss capital. A searching for weblogs of member clubs like these days are helped by the Paris in Pro and Paris Pro initial electronic search process. The search has also surfaced weblogs for the establishment. A fact that the organization helped wish my club a very success, a representative added. In warned a crowd of retirees on September 11th using beer cans in Paris could be a solution for soldiers and retirees, it has been suggested.
Text obtained with 88 NFEs distilled by DiDi-Instruct (3/3). Perplexity=$62.24$, Entropy=$5.17$. The company to extend a ban on bicycles. The entrepreneurs of e-watch-day-lab imposed a ban on passengers from watching their bicycles. Under the ban, the company would be extended into the company. The entrepreneurs have teamed up after the airline imposed a ban. The company ban has already sparked a discussion and debate with passengers, advocacy groups say. The bicycle company ban has already sparked a discussion and debate in India. The entrepreneurs of the watch-gating company have teamed up after the airline imposed the ban. A mistake that turned back decades of the United States presidency has turned back four out of five recordings. The Associated Press reported Thursday that the recordings were meant for a long-ago conversation. The governors campaign vowed never to help his son improve his memory. Assaulting a reporter has turned a mistake of history, which turned back four out of five climactic recordings, forcing the President to listen to a prerecorded conversation. For all the math, its true that the Gang-8 conversation was meant to turn back four times. In a potentially inefficient experiment, the governors campaign supporters are taking a new substance to help improve his poor memory. The problems are with the judiciary and civil courts. The substance is a potential experiment from 1997. He has never played a role. The Industry Association of America will set an industry-wide minimum every year in an effort to help companies take advantage of the problem. Should a company spend more than \$1000 in a year to take advantage, this minimum isnt needed for them. The Utah E-Al Bears have become a favorite of fans. They have a record pace of up to 98 percent temperatures, which translates to degrees Celsius faster than the fall football season. The E-Al have become a staple of bumps, bruises and bumps throughout this season. The computer tells you that your schools sponsors arent as long as your e-alumni. Dependents alum sponsors arent as long as it takes you. There is no trace your alum put on. Now you get a whole e-mail that tells you this.

### v3

@@ E.3 Results on zero-shot perplexities
E.4 Downstream Task Evaluation
We conduct three downstream evaluations used to verify whether DiDi-Instruct preserves the task capability of the teacher’s semantic representations while enabling few-step generation. All experiments use the same 169M architecture unless otherwise specified.
[…]
Table 9: Domain adaptation on MMLU dataset.
[…]
MDLM 9.879 2.814 7.065 20.6% 25.3% 4.7%
SDTT 9.773 2.844 6.929 21.2% 24.3% 3.1%
DiDi-Instruct (Ours) 9.960 2.815 7.145 21.6% 25.1% 3.5%
Scientific domain adaptation. We further evaluate domain transfer to scientific text by conducting 5,000 SFT steps on the PubMed corpus. All models are fine-tuned using AdamW with a learning rate of $5\times 10^{-6}$ and a constant schedule with 1,000 warmup steps. Training uses a per-GPU batch size of 36 with gradient accumulation (effective batch size 72). As reported in Table 10, DiDi-Instruct significantly outperforms the SDTT and DUO baselines. Notably, it achieves a final perplexity of 24.894, which is practically indistinguishable from the teacher’s perplexity, which validates its effectiveness in specialized domains.
Table 10: Domain adaptation on PubMed dataset.
Models NLL (pre) PPL (pre) NLL (post) $\downarrow$ PPL (post) $\downarrow$
MDLM 3.761 42.983 3.217 24.944
SDTT 3.938 51.336 3.243 25.602
DUO 4.028 56.174 3.338 27.734
DiDi-Instruct (Ours ) 3.879 47.563 3.215 24.894
Frozen feature extractor evaluation. To assess the quality of the learned latent representations, we evaluate the models on the GLUE MRPC task (Dolan and Brockett, 2005). The pretrained backbones (MDLM, SDTT, DiDi-Instruct) are kept fixed, and a lightweight classification head is added on top of the encoder’s pooled hidden representation. The head consists of a single linear layer mapping the pooled vector to two logits, and we additionally unfreeze the final encoder block to allow light adaptation. Training uses AdamW with a learning rate of $2\times 10^{-5}$, weight decay 0.01, a batch size of 8, and 10,000 training steps. We report Accuracy and F1 on the validation split.
Results in Table 11 show that DiDi-Instruct attains the highest Accuracy and F1 scores. This confirms that DiDi-Instruct preserves and potentially improves semantic discriminability.
Table 11: Frozen feature evaluation on GLUE MRPC dataset.
Models Accuracy $\uparrow$ F1 Score $\uparrow$
MDLM 59.6% 72.2%
SDTT 59.1% 72.0%
DiDi-Instruct (Ours) 62.3% 73.3%
Across all downstream tasks, DiDi-Instruct matches or exceeds the teacher’s performance, establishing that few-step distillation maintains strong semantic representation quality while delivering significant inference speedups.
E.5 Implementation details of the ablation studies
@@
The baseline model is a one-step generator, mapping a fully masked input $\mathbf{z}_{t}$ ($t=1$) directly to the final output $\mathbf{x}$. With this technique, we decompose the generation into a two-step process $\mathbf{z}_{t}\rightarrow\mathbf{z}_{\tau}\rightarrow\mathbf{x}$ where $t=1$ and $\tau\in(0,1)$. The model first generates an intermediate state $\mathbf{z}_{\tau}$ at a randomly sampled time following $\pi(t)$ and then generates the final output $\mathbf{x}$ from $\mathbf{z}_{\tau}$. The ablation (‘w/o Score Decompose’) reverts to the direct, single-step generation.
Coupled time $t$.
@@
This component synchronizes the time steps used for student generation and discriminator evaluation. In the standard setup, the intermediate denoising time $\tau$ for the student’s two-step generation ($\mathbf{z}_{t}\rightarrow\mathbf{z}_{\tau}$) and the corruption time $\tau^{\prime}$ for preparing the discriminator’s input ($\mathbf{x}\rightarrow\mathbf{z}_{\tau^{\prime}}$ for corrupted student generations, and $\mathbf{x}^{\prime}\rightarrow\mathbf{z}_{\tau^{\prime}}^{\prime}$ for corrupted teacher generations) are coupled, i.e., $\tau=\tau^{\prime}$. The ablation (‘w/o Coupled Time $t$’) decouples them by sampling $\tau$ and $\tau^{\prime}$ independently.
Weight function $\omega(t)$ correction.
@@
The gradient of the objective (5) is weighted by a time-dependent factor $\omega(t)$. The ablation (‘w/o $\omega(t)$ Correction’) uses a constant weight, i.e., $\omega(t)=1$. Our full model applies a theoretically-motivated correction based on the noise schedule’s signal rate $\alpha_{t}$, setting $\omega(t)\propto-\alpha^{\prime}_{t}/(1-\alpha_{t})$, which re-weights the objective based on the difficulty of the denoising step at time $t$.
Time scheduler $\pi(\cdot)$ weighting.
@@
This controls how $\tau$ is sampled and (optionally) whether we apply importance weighting, which also refers to the sampling distribution $\pi(\tau)$ for the intermediate time step $\tau$. The ablation setting (‘w/o $\pi(\tau)$ Weighting’) samples $\tau$ from a uniform distribution. Our method employs a non-uniform Beta distribution as $\pi(\tau)$ to focus the student’s training on specific, more informative intervals of the denoising trajectory. Specifically, we consider heavy-early ($\text{Beta}(2,5)$), and heavy-late ($\text{Beta}(5,2)$) schedules to improve sample quality. Practically, for smaller NFEs (e.g., 8 and 16 NFEs) we bias toward earlier times (e.g., $\text{Beta}(2,5)$); for larger NFEs (e.g., 32, 64, and 128 NFEs) we bias toward later times (e.g., $\text{Beta}(5,2)$).
Regularization.
@@
To stabilize training and maintain generalization, we introduce two regularization terms to the student’s loss. The first is a KL divergence loss, which aligns the student’s output logits with those of the frozen teacher at both steps of the generation process ($\mathbf{z}_{t}\rightarrow\mathbf{z}_{\tau}$ and $\mathbf{z}_{\tau}\rightarrow\mathbf{x}$, where $t=1$ and $\tau\in(0,1)$). Among Forward/Backward/Jeffreys implementations, we found forward KL consistently yielded the most stable optimization and best validation in our cases, so we adopt it for all ablations and apply it with a weight of $0.05$ in the student objective. The second term is an entropy bonus on the student’s output distribution to encourage exploration, which is weighted by $0.0005$. The ablation (‘w/o Regularization’) removes both the KL-divergence and entropy terms from the student’s objective.
Guided inference.
@@
As elaborated in Section 3.4, this technique is applied only during sampling to leverage the trained discriminator and improve sample quality. We employ a hybrid strategy that divides the denoising process into two phases. For the first 50% of NFEs, we use gradient tilting ($M=1$), where the guidance scale $h$ is linearly increased from 30.0 to 40.0 to steer the generation. For the remaining 50% of NFEs, we switch to multi-candidate re-ranking ($h=0$), sampling $M=4$ candidates at each step and selecting the one with the highest reward. The ablation (‘w/o Guided Inference’) disables this process ($h=0,M=1$), reverting to standard unguided ancestral sampling.
E.6 Ablation study result analysis.
@@
E.7 Sensitivity Analysis of RGAS
Figure 3: Sensitivity analysis of RGAS hyperparameters at 16 NFEs. We visualize the landscape of PPL (left) and Sequence Entropy (right) across varying gradient scales $h$, and candidate counts $M$. The heatmap indicates that increasing both $h$ and $M$ consistently reduces PPL, improving generation fidelity. However, larger gradient scales tend to lower sequence entropy, indicating a trade-off between sample quality and diversity.
To understand the interaction between the tilting scale $h$ and candidate number $M$ in RGAS, we conduct a comprehensive grid search to visualize the performance landscape, and a global sensitivity analysis using Functional ANOVA (fANOVA) (Hutter et al., 2014) to quantify hyperparameter importance.
Performance landscape analysis. We first study the joint effect of the RGAS hyperparameters $h$ and $M$ on generation quality. At a fixed budget of 16 NFEs and batch size 16, we perform a dense grid search, sweeping $h\in[0,100]$ and $M\in\{1,2,4,8,16,32\}$. Figure 3 presents the resulting PPL and Entropy. The resulting landscapes show a broad valley: increasing $h$ from 0 to 100 consistently reduces PPL, while entropy only gradually decreases and remains close to the teacher’s entropy. This indicates that RGAS is relatively robust across a wide range of $(h,M)$ values and that there is no narrow “knife-edge” region required for good performance.
Global sensitivity with fANOVA. To formally quantify the relative importance of $h$ and $M$, we conduct a global sensitivity analysis using Functional ANOVA. We adopt a Latin Hypercube Sampling strategy and generate 1,000 hyperparameter pairs, with $h$ sampled continuously from $[20,60]$ (a range that balances PPL and entropy) and $M$ sampled as an integer from $[1,32]$. For each configuration, we run RGAS with 16 generated sequences and record the resulting PPL. This experiment is repeated separately for NFEs =8,16,32=8,16,32, and the resulting performance is decomposed into variance contributions from $h$ and $M$.
Table 12: Variance decomposition of PPL via fANOVA.
Parameter 8 NFEs 16 NFEs 32 NFEs
Gradient Scale ($h$) 92.7% 90.0% 63.7%
Candidate Count ($M$) 7.3% 6.0% 36.3%
Table 12 summarizes the fANOVA results. At 8 and 16 NFEs, $h$ explains around $90\%$ of the variance in PPL, while $M$ accounts for less than $10\%$. At 32 NFEs, the importance of $M$ increases, but $h$ still contributes the majority of the variance. These findings suggest that: (i) in the low-NFE regime, $h$ is the dominant hyperparameter and should be prioritized during tuning; and (ii) as the sampling budget increases, $M$ becomes a more relevant knob for refining performance.
E.8 Model Scaling Up
@@
E.9 Protein Sequence Generation
Figure 5: pLDDT comparison between DiDi-Instruct and the DPLM-150M across different sequence lengths ($L=100,200,300,400,500$) and numbers of function evaluations (NFEs). Our method consistently outperforms the teacher, achieving up to +10 pLDDT gains at shorter sequence lengths (e.g., $L=100$) and maintaining superior structural confidence across all lengths, even with substantially fewer sampling steps. Moreover, the distilled student exhibits more stable performance across NFEs, whereas the teacher shows larger variability as the number of steps increases.
Dataset and evaluation metric. The UniRef50 dataset (Suzek et al., 2014) contains approximately 45 million protein sequences, comprising roughly 14 billion amino acid tokens. Following previous work (Wang et al., 2024), we adopt a vocabulary of 33 tokens and a maximum sequence length of 1024, with longer proteins chunked into shorter chunks. We compute the predicted local distance difference test (pLDDT) score using the ESMFold model (Lin et al., 2023) to assess the foldability of the generated protein sequences, with values above 70 indicating high structural confidence. We report the pLDDT score across different sequence lengths and varying NFEs.
@@ E.9 Protein Sequence Generation
Figure 7: Visualization examples of high-confidence protein sequences ($\text{pLDDT}>70$) generated by the DiDi-Instruct model, with lengths ranging from 50 to 500 across different numbers of function evaluations (NFEs).
Sequence diversity analysis. High predicted structural confidence (pLDDT) can sometimes indicate a risk of mode collapse, where the model generates repetitive samples. Therefore, we evaluate sequence-level diversity using MMseqs2 clustering (Steinegger and Söding, 2017) with a 30% sequence identity and 80% coverage threshold—standard criteria for determining biologically meaningful variation. Across low-NFE settings (32–64), DiDi-Instruct exhibits diversity comparable to DPLM, which reflects the stochastic nature of early denoising. As the number of sampling steps increases (128–512 NFEs), DiDi-Instruct yields slightly lower entropy and moderately larger cluster sizes, indicating a more concentrated output distribution. Importantly, diversity levels remain competitive with DPLM across all settings, and no regime shows abnormal cluster growth or a collapse of entropy. Noted that genuine mode collapse in this setting would correspond to near-zero cluster entropy together with an average cluster size approaching the full sample count, i.e., all sequences falling into a single MMseqs2 cluster. This collapse behavior is not observed in any configuration.
This behavior is consistent with our text-generation ablations, where diversity naturally decreases as NFEs grow and the denoising trajectory becomes more deterministic. Moreover, DiDi-Instruct is explicitly designed for few-step generation; mild diversity reduction at very large NFEs does not contradict our intended operating regime. Overall, these results confirm that DiDi-Instruct maintains biologically meaningful variability across the generation budget while avoiding collapse to a small subset of high-pLDDT sequences.
Table 13: Protein sequence diversity evaluation using MMseqs2 clustering. We compare the Average Cluster Size and Cluster Entropy across different NFEs.
NFEs Methods Cluster Entropy $\uparrow$ Average Cluster Size $\downarrow$
100 200 300 400 500
32 DPLM 3.68 3.68 3.55 3.65 3.68 1.02
DiDi-Instruct (Ours ) 3.68 3.52 3.15 3.58 3.55 1.11
64 DPLM 3.68 3.51 3.45 3.68 3.53 1.07
DiDi-Instruct (Ours) 3.68 3.65 3.45 3.45 3.54 1.09
128 DPLM 3.68 3.65 3.50 3.52 3.61 1.05
DiDi-Instruct (Ours) 3.62 3.20 3.42 3.45 3.38 1.21
256 DPLM 3.61 3.42 3.41 3.19 3.44 1.19
DiDi-Instruct (Ours) 3.61 3.11 3.41 2.87 3.07 1.40
512 DPLM 3.58 3.23 3.02 3.11 2.97 1.41
DiDi-Instruct (Ours) 3.54 3.11 2.95 2.77 2.79 1.53
E.10 Text examples
@@
Text obtained with $1024$ NFEs teacher model. Perplexity=$40.48$, Entropy=$5.22$. He was mid-morning early, and I started forming a motorcade for him. He stopped every NYPD car, and I radioed back to every bicycle officer that was in there, to protect him. I did not hear you said to be involved in it. That you were involved with that campaign. I never asked anybody to test it. You did need help you leaked it, not made it for yourself. He parked and it was then that I saw that president and ended up there, speaking with a journalist. The Prince was granted anonymity and noted that the FBI was aware of the familiarity with Russia and was conducting the investigation as appropriate. In some accounts, Casey Wilson, senior counsel for the Bonden Bisson, told The Post that the President is aware of the inquiry. But Mr. Bakers testimony is the first time someone has long taken something of concern into the hands of their officials probing Trumps personal relationships with foreign officials. Former FBI regional director James Kogan, a former officer who met with President Trump, was once briefed about the inquiry from the Russian District Attorney General. Victor Atkinson, an FBI chief-officer who is retired, hacked the personal PINs and cell phone lines of the two Trump officials to conceal the information, as shown inside the countrys electronic networks. Anonymous messages can appear in our apps like Twitter and Facebook and if ones political actions resists they become blocked.
@@
Text obtained with 88 NFEs distilled by DiDi-Instruct. Perplexity=$62.24$, Entropy=$5.17$. An engineers run up against a system of dynamic memory-producing technology. An examiners mechanism of memory has to reduce property crimes against the technology. Indias memory system has run into a deficit of memory and that economic damage in India has run up against a system of Now and Then. The new system will be designed as an economically secure and efficient system in determining how residents receive support. A loss of support from memory management systems wont go too far from promise. With a deficit of 1 million for residents, the Navajo Nation is testing a new system that wont prevent you from receiving your financial stream. The Navajo Nation faces a deficit of 1.88 million in systems that are designed to enhance what residents receive. Its a potential loss of support from a system that has been receiving both positive and negative feedback from the community. A presidential candidate who once served during four tours in Iraq and Afghanistan refers to his new record as a figurative use of pop culture. I got a disappointed order from the State Department, the new record is to go against the American health care system.